# Nondeterministic Polynomial-time Problem Challenge: An Ever-Scaling Reasoning Benchmark for LLMs

### v1

@@ Appendix G Full Results over Problems
QwQ-32B $1.00_{1.00}^{1.00}$ $0.92_{0.87}^{0.97}$ $0.28_{0.20}^{0.33}$ $0.19_{0.13}^{0.23}$ $0.17_{0.10}^{0.23}$ $0.06_{0.00}^{0.13}$ $0.08_{0.03}^{0.13}$ $0.01_{0.00}^{0.03}$ $0.00_{0.00}^{0.00}$ $0.00_{0.00}^{0.00}$
DeepSeek-R1-32B $0.58_{0.33}^{0.70}$ $0.24_{0.10}^{0.40}$ $0.16_{0.07}^{0.23}$ $0.12_{0.07}^{0.23}$ $0.10_{0.07}^{0.17}$ $0.03_{0.03}^{0.03}$ $0.02_{0.00}^{0.03}$ $0.00_{0.00}^{0.00}$ $0.00_{0.00}^{0.00}$ $0.00_{0.00}^{0.00}$
GPT-4o-mini $0.32_{0.20}^{0.47}$ $0.08_{0.03}^{0.13}$ $0.02_{0.00}^{0.03}$ $0.01_{0.00}^{0.03}$ $0.00_{0.00}^{0.00}$ $0.01_{0.00}^{0.03}$ $0.00_{0.00}^{0.00}$ $0.01_{0.00}^{0.03}$ $0.00_{0.00}^{0.00}$ $0.00_{0.00}^{0.00}$
GPT-4o $0.81_{0.77}^{0.83}$ $0.47_{0.30}^{0.57}$ $0.11_{0.07}^{0.17}$ $0.10_{0.03}^{0.17}$ $0.06_{0.03}^{0.10}$ $0.16_{0.03}^{0.27}$ $0.06_{0.03}^{0.10}$ $0.12_{0.10}^{0.13}$ $0.07_{0.03}^{0.10}$ $0.03_{0.00}^{0.07}$
Claude-3.7-Sonnet $0.99_{0.97}^{1.00}$ $0.97_{0.93}^{1.00}$ $0.78_{0.70}^{0.90}$ $0.51_{0.47}^{0.57}$ $0.68_{0.60}^{0.77}$ $0.74_{0.70}^{0.80}$ $0.77_{0.73}^{0.80}$ $0.88_{0.83}^{0.93}$ $0.82_{0.77}^{0.90}$ $0.74_{0.70}^{0.80}$
DeepSeek-V3 $0.80_{0.73}^{0.83}$ $0.52_{0.37}^{0.67}$ $0.49_{0.43}^{0.53}$ $0.46_{0.40}^{0.53}$ $0.44_{0.40}^{0.53}$ $0.40_{0.30}^{0.57}$ $0.24_{0.13}^{0.37}$ $0.22_{0.17}^{0.27}$ $0.08_{0.03}^{0.10}$ $0.02_{0.00}^{0.03}$
@@
QwQ-32B $0.72_{0.30}^{1.00}$ $0.56_{0.17}^{0.80}$ $0.19_{0.07}^{0.27}$ $0.16_{0.07}^{0.23}$ $0.03_{0.03}^{0.03}$ $0.00_{0.00}^{0.00}$ $0.00_{0.00}^{0.00}$ $0.00_{0.00}^{0.00}$ $0.00_{0.00}^{0.00}$ $0.00_{0.00}^{0.00}$
DeepSeek-R1-32B $0.84_{0.77}^{0.90}$ $0.62_{0.50}^{0.70}$ $0.11_{0.10}^{0.13}$ $0.08_{0.07}^{0.10}$ $0.00_{0.00}^{0.00}$ $0.00_{0.00}^{0.00}$ $0.00_{0.00}^{0.00}$ $0.00_{0.00}^{0.00}$ $0.00_{0.00}^{0.00}$ $0.00_{0.00}^{0.00}$
GPT-4o-mini $0.49_{0.43}^{0.53}$ $0.23_{0.17}^{0.27}$ $0.03_{0.00}^{0.07}$ $0.00_{0.00}^{0.00}$ $0.00_{0.00}^{0.00}$ $0.00_{0.00}^{0.00}$ $0.00_{0.00}^{0.00}$ $0.00_{0.00}^{0.00}$ $0.00_{0.00}^{0.00}$ $0.00_{0.00}^{0.00}$
GPT-4o $0.67_{0.60}^{0.70}$ $0.43_{0.33}^{0.57}$ $0.08_{0.07}^{0.10}$ $0.03_{0.03}^{0.03}$ $0.01_{0.00}^{0.03}$ $0.00_{0.00}^{0.00}$ $0.00_{0.00}^{0.00}$ $0.00_{0.00}^{0.00}$ $0.00_{0.00}^{0.00}$ $0.00_{0.00}^{0.00}$
Claude-3.7-Sonnet $0.96_{0.87}^{1.00}$ $0.97_{0.97}^{0.97}$ $0.78_{0.77}^{0.80}$ $0.59_{0.47}^{0.67}$ $0.10_{0.07}^{0.13}$ $0.00_{0.00}^{0.00}$ $0.00_{0.00}^{0.00}$ $0.00_{0.00}^{0.00}$ $0.00_{0.00}^{0.00}$ $0.00_{0.00}^{0.00}$
DeepSeek-V3 $0.97_{0.97}^{0.97}$ $0.89_{0.83}^{0.93}$ $0.38_{0.37}^{0.40}$ $0.19_{0.10}^{0.30}$ $0.04_{0.03}^{0.07}$ $0.02_{0.00}^{0.03}$ $0.00_{0.00}^{0.00}$ $0.00_{0.00}^{0.00}$ $0.00_{0.00}^{0.00}$ $0.00_{0.00}^{0.00}$
DeepSeek-V3-2503 $1.00_{1.00}^{1.00}$ $1.00_{1.00}^{1.00}$ $0.68_{0.63}^{0.70}$ $0.64_{0.57}^{0.73}$ $0.30_{0.20}^{0.37}$ $0.17_{0.13}^{0.20}$ $0.08_{0.00}^{0.13}$ $0.01_{0.00}^{0.03}$ $0.08_{0.00}^{0.13}$ $0.00_{0.00}^{0.00}$
DeepSeek-R1 $1.00_{1.00}^{1.00}$ $1.00_{1.00}^{1.00}$ $1.00_{1.00}^{1.00}$ $0.97_{0.93}^{1.00}$ $0.82_{0.77}^{0.90}$ $0.68_{0.63}^{0.73}$ $0.27_{0.20}^{0.33}$ $0.17_{0.13}^{0.20}$ $0.09_{0.07}^{0.13}$ $0.03_{0.00}^{0.07}$
o1-mini $0.57_{0.50}^{0.70}$ $0.59_{0.50}^{0.63}$ $0.44_{0.33}^{0.63}$ $0.46_{0.43}^{0.50}$ $0.11_{0.07}^{0.17}$ $0.03_{0.00}^{0.07}$ $0.00_{0.00}^{0.00}$ $0.00_{0.00}^{0.00}$ $0.00_{0.00}^{0.00}$ $0.00_{0.00}^{0.00}$
o3-mini $0.94_{0.90}^{0.97}$ $0.99_{0.97}^{1.00}$ $0.94_{0.93}^{0.97}$ $0.96_{0.90}^{1.00}$ $0.81_{0.77}^{0.87}$ $0.66_{0.53}^{0.77}$ $0.30_{0.20}^{0.43}$ $0.27_{0.20}^{0.30}$ $0.27_{0.23}^{0.30}$ $0.13_{0.10}^{0.17}$
@@
QwQ-32B $1.00_{1.00}^{1.00}$ $0.98_{0.97}^{1.00}$ $0.93_{0.90}^{0.97}$ $0.94_{0.93}^{0.97}$ $0.33_{0.07}^{0.83}$ $0.06_{0.00}^{0.10}$ $0.00_{0.00}^{0.00}$ $0.00_{0.00}^{0.00}$ $0.00_{0.00}^{0.00}$ $0.00_{0.00}^{0.00}$
DeepSeek-R1-32B $0.87_{0.77}^{1.00}$ $0.42_{0.23}^{0.57}$ $0.09_{0.03}^{0.13}$ $0.00_{0.00}^{0.00}$ $0.00_{0.00}^{0.00}$ $0.01_{0.00}^{0.03}$ $0.00_{0.00}^{0.00}$ $0.00_{0.00}^{0.00}$ $0.00_{0.00}^{0.00}$ $0.00_{0.00}^{0.00}$
GPT-4o-mini $0.43_{0.27}^{0.57}$ $0.09_{0.07}^{0.10}$ $0.02_{0.00}^{0.03}$ $0.00_{0.00}^{0.00}$ $0.00_{0.00}^{0.00}$ $0.00_{0.00}^{0.00}$ $0.00_{0.00}^{0.00}$ $0.00_{0.00}^{0.00}$ $0.00_{0.00}^{0.00}$ $0.00_{0.00}^{0.00}$
GPT-4o $0.64_{0.53}^{0.83}$ $0.24_{0.17}^{0.37}$ $0.13_{0.03}^{0.20}$ $0.10_{0.07}^{0.13}$ $0.02_{0.00}^{0.07}$ $0.00_{0.00}^{0.00}$ $0.00_{0.00}^{0.00}$ $0.00_{0.00}^{0.00}$ $0.00_{0.00}^{0.00}$ $0.00_{0.00}^{0.00}$
Claude-3.7-Sonnet $0.96_{0.93}^{0.97}$ $0.84_{0.80}^{0.90}$ $0.76_{0.67}^{0.80}$ $0.59_{0.43}^{0.70}$ $0.21_{0.13}^{0.33}$ $0.09_{0.07}^{0.10}$ $0.07_{0.03}^{0.10}$ $0.00_{0.00}^{0.00}$ $0.00_{0.00}^{0.00}$ $0.00_{0.00}^{0.00}$
DeepSeek-V3 $0.74_{0.57}^{0.83}$ $0.32_{0.23}^{0.47}$ $0.08_{0.00}^{0.13}$ $0.03_{0.00}^{0.10}$ $0.00_{0.00}^{0.00}$ $0.00_{0.00}^{0.00}$ $0.00_{0.00}^{0.00}$ $0.00_{0.00}^{0.00}$ $0.00_{0.00}^{0.00}$ $0.00_{0.00}^{0.00}$
DeepSeek-V3-2503 $0.94_{0.93}^{0.97}$ $0.76_{0.70}^{0.87}$ $0.49_{0.43}^{0.60}$ $0.31_{0.10}^{0.47}$ $0.07_{0.03}^{0.10}$ $0.03_{0.00}^{0.07}$ $0.01_{0.00}^{0.03}$ $0.00_{0.00}^{0.00}$ $0.00_{0.00}^{0.00}$ $0.00_{0.00}^{0.00}$
DeepSeek-R1 $1.00_{1.00}^{1.00}$ $1.00_{1.00}^{1.00}$ $0.98_{0.97}^{1.00}$ $0.97_{0.93}^{1.00}$ $0.93_{0.90}^{0.97}$ $0.91_{0.83}^{0.97}$ $0.91_{0.87}^{0.97}$ $0.57_{0.50}^{0.67}$ $0.27_{0.17}^{0.37}$ $0.02_{0.00}^{0.03}$
o1-mini $0.87_{0.83}^{0.93}$ $0.89_{0.87}^{0.90}$ $0.81_{0.73}^{0.87}$ $0.77_{0.73}^{0.83}$ $0.38_{0.30}^{0.47}$ $0.26_{0.23}^{0.27}$ $0.11_{0.07}^{0.20}$ $0.01_{0.00}^{0.03}$ $0.00_{0.00}^{0.00}$ $0.00_{0.00}^{0.00}$
@@
QwQ-32B $0.61_{0.50}^{0.80}$ $0.41_{0.30}^{0.53}$ $0.42_{0.30}^{0.53}$ $0.56_{0.50}^{0.60}$ $0.26_{0.23}^{0.30}$ $0.19_{0.13}^{0.27}$ $0.02_{0.00}^{0.07}$ $0.00_{0.00}^{0.00}$ $0.00_{0.00}^{0.00}$ $0.00_{0.00}^{0.00}$
DeepSeek-R1-32B $0.88_{0.87}^{0.90}$ $0.62_{0.53}^{0.73}$ $0.30_{0.23}^{0.40}$ $0.13_{0.03}^{0.20}$ $0.02_{0.00}^{0.03}$ $0.01_{0.00}^{0.03}$ $0.00_{0.00}^{0.00}$ $0.00_{0.00}^{0.00}$ $0.00_{0.00}^{0.00}$ $0.00_{0.00}^{0.00}$
GPT-4o-mini $0.93_{0.90}^{0.96}$ $0.34_{0.27}^{0.40}$ $0.12_{0.07}^{0.20}$ $0.07_{0.00}^{0.10}$ $0.00_{0.00}^{0.00}$ $0.00_{0.00}^{0.00}$ $0.00_{0.00}^{0.00}$ $0.00_{0.00}^{0.00}$ $0.00_{0.00}^{0.00}$ $0.00_{0.00}^{0.00}$
GPT-4o $0.97_{0.97}^{0.97}$ $0.76_{0.73}^{0.80}$ $0.59_{0.50}^{0.67}$ $0.40_{0.33}^{0.47}$ $0.22_{0.10}^{0.33}$ $0.16_{0.03}^{0.23}$ $0.08_{0.07}^{0.10}$ $0.02_{0.00}^{0.07}$ $0.00_{0.00}^{0.00}$ $0.00_{0.00}^{0.00}$
Claude-3.7-Sonnet $1.00_{1.00}^{1.00}$ $0.98_{0.97}^{1.00}$ $0.90_{0.83}^{0.93}$ $0.83_{0.77}^{0.90}$ $0.86_{0.80}^{0.90}$ $0.80_{0.73}^{0.83}$ $0.54_{0.47}^{0.70}$ $0.51_{0.50}^{0.53}$ $0.08_{0.03}^{0.10}$ $0.06_{0.00}^{0.10}$
DeepSeek-V3 $0.98_{0.93}^{1.00}$ $0.90_{0.90}^{0.90}$ $0.74_{0.60}^{0.83}$ $0.62_{0.50}^{0.77}$ $0.49_{0.33}^{0.67}$ $0.49_{0.33}^{0.73}$ $0.17_{0.13}^{0.23}$ $0.07_{0.00}^{0.13}$ $0.02_{0.00}^{0.03}$ $0.00_{0.00}^{0.00}$
DeepSeek-V3-2503 $1.00_{1.00}^{1.00}$ $0.94_{0.90}^{0.97}$ $0.96_{0.87}^{1.00}$ $0.83_{0.80}^{0.87}$ $0.70_{0.67}^{0.77}$ $0.66_{0.57}^{0.70}$ $0.39_{0.27}^{0.47}$ $0.10_{0.10}^{0.10}$ $0.01_{0.00}^{0.03}$ $0.01_{0.00}^{0.03}$
DeepSeek-R1 $1.00_{1.00}^{1.00}$ $0.99_{0.97}^{1.00}$ $0.97_{0.97}^{0.97}$ $0.99_{0.97}^{1.00}$ $0.87_{0.87}^{0.87}$ $0.78_{0.77}^{0.80}$ $0.62_{0.57}^{0.67}$ $0.24_{0.20}^{0.30}$ $0.03_{0.00}^{0.10}$ $0.00_{0.00}^{0.00}$
o1-mini $0.84_{0.73}^{0.90}$ $0.89_{0.87}^{0.93}$ $0.67_{0.60}^{0.77}$ $0.57_{0.43}^{0.63}$ $0.34_{0.23}^{0.47}$ $0.37_{0.23}^{0.43}$ $0.18_{0.07}^{0.30}$ $0.01_{0.00}^{0.03}$ $0.00_{0.00}^{0.00}$ $0.00_{0.00}^{0.00}$
o3-mini $0.79_{0.73}^{0.87}$ $0.62_{0.53}^{0.67}$ $0.53_{0.47}^{0.63}$ $0.28_{0.20}^{0.33}$ $0.31_{0.23}^{0.37}$ $0.30_{0.17}^{0.47}$ $0.30_{0.20}^{0.37}$ $0.19_{0.17}^{0.20}$ $0.12_{0.07}^{0.17}$ $0.07_{0.00}^{0.13}$

### v2

@@ Appendix G Full Results over Problems
QwQ-32B $1.00_{1.00}^{1.00}$ $0.92_{0.87}^{0.97}$ $0.28_{0.20}^{0.33}$ $0.19_{0.13}^{0.23}$ $17_{0.10}^{0.23} 0.0$$6_{0.00}^{0.13} 0.0$$8_{0.03}^{0.13} 0.0$$1_{0.00}^{0.03} 0.0$$_{0.00}^{0.00} 0.000$$0_{0.00}^{0.00}$
DeepSeek-R1-32B $0.58_{0.33}^{0.70}$ $0.24_{0.10}^{0.40}$ $0.16_{0.07}^{0.23}$ $0.12_{0.07}^{0.23}$ $10_{0.07}^{0.17} 0.0$$3_{0.03}^{0.03} 0.0$$2_{0.00}^{0.03} 0.0$$_{0.00}^{0.00} 0.000$$_{0.00}^{0.00} 0.0$$00_{0.00}^{0.00}$
GPT-4o-mini $32_{0.20}^{0.47} 0.0$$8_{0.03}^{0.13} 0.0$$2_{0.00}^{0.03} 0.0$$1_{0.00}^{0.03} 0.0$$0_{0.00}^{0.00} 0.0$$1_{0.00}^{0.03} 0.0$$0_{0.00}^{0.00} 0.0$$1_{0.00}^{0.03} 0.0$$_{0.00}^{0.00} 0.000$$0_{0.00}^{0.00}$
GPT-4o $0.81_{0.77}^{0.83}$ $0.47_{0.30}^{0.57}$ $0.11_{0.07}^{0.17}$ $10_{0.03}^{0.17} 0.0$$06_{0.03}^{0.10}$ $16_{0.03}^{0.27} 0.0$$06_{0.03}^{0.10}$ $12_{0.10}^{0.13} 0.0$$7_{0.03}^{0.10} 0.0$$03_{0.00}^{0.07}$
Claude-3.7-Sonnet $0.99_{0.97}^{1.00}$ $0.97_{0.93}^{1.00}$ $0.78_{0.70}^{0.90}$ $0.51_{0.47}^{0.57}$ $0.68_{0.60}^{0.77}$ $4_{0.70}^{0.80} 0.770$$7_{0.73}^{0.80}$ $0.88_{0.83}^{0.93}$ $0.82_{0.77}^{0.90}$ $0.74_{0.70}^{0.80}$
DeepSeek-V3 $0.80_{0.73}^{0.83}$ $0.52_{0.37}^{0.67}$ $0.49_{0.43}^{0.53}$ $6_{0.40}^{0.53} 0.440$$4_{0.40}^{0.53}$ $0.40_{0.30}^{0.57}$ $4_{0.13}^{0.37} 0.220$$2_{0.17}^{0.27}$ $08_{0.03}^{0.10} 0.0$$02_{0.00}^{0.03}$
@@
QwQ-32B $0.72_{0.30}^{1.00}$ $0.56_{0.17}^{0.80}$ $0.19_{0.07}^{0.27}$ $16_{0.07}^{0.23} 0.0$$3_{0.03}^{0.03} 0.0$$_{0.00}^{0.00} 0.000$$_{0.00}^{0.00} 0.0$$_{0.00}^{0.00} 0.000$$_{0.00}^{0.00} 0.0$$00_{0.00}^{0.00}$
DeepSeek-R1-32B $0.84_{0.77}^{0.90}$ $0.62_{0.50}^{0.70}$ $11_{0.10}^{0.13} 0.0$$8_{0.07}^{0.10} 0.0$$_{0.00}^{0.00} 0.000$$_{0.00}^{0.00} 0.0$$_{0.00}^{0.00} 0.000$$_{0.00}^{0.00} 0.0$$_{0.00}^{0.00} 0.000$$0_{0.00}^{0.00}$
GPT-4o-mini $0.49_{0.43}^{0.53}$ $23_{0.17}^{0.27} 0.0$$3_{0.00}^{0.07} 0.0$$_{0.00}^{0.00} 0.000$$_{0.00}^{0.00} 0.0$$_{0.00}^{0.00} 0.000$$_{0.00}^{0.00} 0.0$$_{0.00}^{0.00} 0.000$$_{0.00}^{0.00} 0.0$$00_{0.00}^{0.00}$
GPT-4o $0.67_{0.60}^{0.70}$ $43_{0.33}^{0.57} 0.0$$8_{0.07}^{0.10} 0.0$$3_{0.03}^{0.03} 0.0$$1_{0.00}^{0.03} 0.0$$_{0.00}^{0.00} 0.000$$_{0.00}^{0.00} 0.0$$_{0.00}^{0.00} 0.000$$_{0.00}^{0.00} 0.0$$00_{0.00}^{0.00}$
Claude-3.7-Sonnet $0.96_{0.87}^{1.00}$ $0.97_{0.97}^{0.97}$ $0.78_{0.77}^{0.80}$ $0.59_{0.47}^{0.67}$ $10_{0.07}^{0.13} 0.0$$_{0.00}^{0.00} 0.000$$_{0.00}^{0.00} 0.0$$_{0.00}^{0.00} 0.000$$_{0.00}^{0.00} 0.0$$00_{0.00}^{0.00}$
DeepSeek-V3 $0.97_{0.97}^{0.97}$ $0.89_{0.83}^{0.93}$ $0.38_{0.37}^{0.40}$ $19_{0.10}^{0.30} 0.0$$4_{0.03}^{0.07} 0.0$$2_{0.00}^{0.03} 0.0$$_{0.00}^{0.00} 0.000$$_{0.00}^{0.00} 0.0$$_{0.00}^{0.00} 0.000$$0_{0.00}^{0.00}$
DeepSeek-V3-2503 $_{1.00}^{1.00} 1.001.00$$_{1.00}^{1.00}$ $0.68_{0.63}^{0.70}$ $0.64_{0.57}^{0.73}$ $0.30_{0.20}^{0.37}$ $17_{0.13}^{0.20} 0.0$$8_{0.00}^{0.13} 0.0$$1_{0.00}^{0.03} 0.0$$8_{0.00}^{0.13} 0.0$$00_{0.00}^{0.00}$
DeepSeek-R1 $_{1.00}^{1.00} 1.001.00$$_{1.00}^{1.00}$ $1.00_{1.00}^{1.00}$ $0.97_{0.93}^{1.00}$ $0.82_{0.77}^{0.90}$ $0.68_{0.63}^{0.73}$ $0.27_{0.20}^{0.33}$ $17_{0.13}^{0.20} 0.0$$9_{0.07}^{0.13} 0.0$$03_{0.00}^{0.07}$
o1-mini $0.57_{0.50}^{0.70}$ $0.59_{0.50}^{0.63}$ $0.44_{0.33}^{0.63}$ $0.46_{0.43}^{0.50}$ $11_{0.07}^{0.17} 0.0$$3_{0.00}^{0.07} 0.0$$_{0.00}^{0.00} 0.000$$_{0.00}^{0.00} 0.0$$_{0.00}^{0.00} 0.000$$0_{0.00}^{0.00}$
o3-mini $4_{0.90}^{0.97} 0.990$$9_{0.97}^{1.00}$ $0.94_{0.93}^{0.97}$ $0.96_{0.90}^{1.00}$ $0.81_{0.77}^{0.87}$ $0.66_{0.53}^{0.77}$ $0.30_{0.20}^{0.43}$ $0.27_{0.20}^{0.30}$ $0.27_{0.23}^{0.30}$ $0.13_{0.10}^{0.17}$
@@
QwQ-32B $1.00_{1.00}^{1.00}$ $0.98_{0.97}^{1.00}$ $0.93_{0.90}^{0.97}$ $0.94_{0.93}^{0.97}$ $33_{0.07}^{0.83} 0.0$$6_{0.00}^{0.10} 0.0$$_{0.00}^{0.00} 0.000$$_{0.00}^{0.00} 0.0$$_{0.00}^{0.00} 0.000$$0_{0.00}^{0.00}$
DeepSeek-R1-32B $0.87_{0.77}^{1.00}$ $42_{0.23}^{0.57} 0.0$$9_{0.03}^{0.13} 0.0$$_{0.00}^{0.00} 0.000$$_{0.00}^{0.00} 0.0$$1_{0.00}^{0.03} 0.0$$_{0.00}^{0.00} 0.000$$_{0.00}^{0.00} 0.0$$_{0.00}^{0.00} 0.000$$0_{0.00}^{0.00}$
GPT-4o-mini $43_{0.27}^{0.57} 0.0$$9_{0.07}^{0.10} 0.0$$2_{0.00}^{0.03} 0.0$$_{0.00}^{0.00} 0.000$$_{0.00}^{0.00} 0.0$$_{0.00}^{0.00} 0.000$$_{0.00}^{0.00} 0.0$$_{0.00}^{0.00} 0.000$$_{0.00}^{0.00} 0.0$$00_{0.00}^{0.00}$
GPT-4o $0.64_{0.53}^{0.83}$ $0.24_{0.17}^{0.37}$ $0.13_{0.03}^{0.20}$ $10_{0.07}^{0.13} 0.0$$2_{0.00}^{0.07} 0.0$$_{0.00}^{0.00} 0.000$$_{0.00}^{0.00} 0.0$$_{0.00}^{0.00} 0.000$$_{0.00}^{0.00} 0.0$$00_{0.00}^{0.00}$
Claude-3.7-Sonnet $0.96_{0.93}^{0.97}$ $0.84_{0.80}^{0.90}$ $0.76_{0.67}^{0.80}$ $0.59_{0.43}^{0.70}$ $21_{0.13}^{0.33} 0.0$$9_{0.07}^{0.10} 0.0$$7_{0.03}^{0.10} 0.0$$_{0.00}^{0.00} 0.000$$_{0.00}^{0.00} 0.0$$00_{0.00}^{0.00}$
DeepSeek-V3 $0.74_{0.57}^{0.83}$ $32_{0.23}^{0.47} 0.0$$8_{0.00}^{0.13} 0.0$$3_{0.00}^{0.10} 0.0$$_{0.00}^{0.00} 0.000$$_{0.00}^{0.00} 0.0$$_{0.00}^{0.00} 0.000$$_{0.00}^{0.00} 0.0$$_{0.00}^{0.00} 0.000$$0_{0.00}^{0.00}$
DeepSeek-V3-2503 $0.94_{0.93}^{0.97}$ $0.76_{0.70}^{0.87}$ $0.49_{0.43}^{0.60}$ $31_{0.10}^{0.47} 0.0$$7_{0.03}^{0.10} 0.0$$3_{0.00}^{0.07} 0.0$$1_{0.00}^{0.03} 0.0$$_{0.00}^{0.00} 0.000$$_{0.00}^{0.00} 0.0$$00_{0.00}^{0.00}$
DeepSeek-R1 $_{1.00}^{1.00} 1.001.00$$_{1.00}^{1.00}$ $0.98_{0.97}^{1.00}$ $0.97_{0.93}^{1.00}$ $0.93_{0.90}^{0.97}$ $0.91_{0.83}^{0.97}$ $0.91_{0.87}^{0.97}$ $0.57_{0.50}^{0.67}$ $27_{0.17}^{0.37} 0.0$$02_{0.00}^{0.03}$
o1-mini $0.87_{0.83}^{0.93}$ $0.89_{0.87}^{0.90}$ $0.81_{0.73}^{0.87}$ $0.77_{0.73}^{0.83}$ $0.38_{0.30}^{0.47}$ $0.26_{0.23}^{0.27}$ $11_{0.07}^{0.20} 0.0$$1_{0.00}^{0.03} 0.0$$_{0.00}^{0.00} 0.000$$0_{0.00}^{0.00}$
@@
QwQ-32B $0.61_{0.50}^{0.80}$ $0.41_{0.30}^{0.53}$ $0.42_{0.30}^{0.53}$ $0.56_{0.50}^{0.60}$ $0.26_{0.23}^{0.30}$ $19_{0.13}^{0.27} 0.0$$2_{0.00}^{0.07} 0.0$$_{0.00}^{0.00} 0.000$$_{0.00}^{0.00} 0.0$$00_{0.00}^{0.00}$
DeepSeek-R1-32B $0.88_{0.87}^{0.90}$ $0.62_{0.53}^{0.73}$ $0.30_{0.23}^{0.40}$ $13_{0.03}^{0.20} 0.0$$2_{0.00}^{0.03} 0.0$$1_{0.00}^{0.03} 0.0$$_{0.00}^{0.00} 0.000$$_{0.00}^{0.00} 0.0$$_{0.00}^{0.00} 0.000$$0_{0.00}^{0.00}$
GPT-4o-mini $0.93_{0.90}^{0.96}$ $0.34_{0.27}^{0.40}$ $12_{0.07}^{0.20} 0.0$$7_{0.00}^{0.10} 0.0$$_{0.00}^{0.00} 0.000$$_{0.00}^{0.00} 0.0$$_{0.00}^{0.00} 0.000$$_{0.00}^{0.00} 0.0$$_{0.00}^{0.00} 0.000$$0_{0.00}^{0.00}$
GPT-4o $0.97_{0.97}^{0.97}$ $0.76_{0.73}^{0.80}$ $0.59_{0.50}^{0.67}$ $0.40_{0.33}^{0.47}$ $0.22_{0.10}^{0.33}$ $16_{0.03}^{0.23} 0.0$$8_{0.07}^{0.10} 0.0$$2_{0.00}^{0.07} 0.0$$_{0.00}^{0.00} 0.000$$0_{0.00}^{0.00}$
Claude-3.7-Sonnet $1.00_{1.00}^{1.00}$ $0.98_{0.97}^{1.00}$ $0.90_{0.83}^{0.93}$ $0.83_{0.77}^{0.90}$ $0.86_{0.80}^{0.90}$ $0.80_{0.73}^{0.83}$ $0.54_{0.47}^{0.70}$ $51_{0.50}^{0.53} 0.0$$8_{0.03}^{0.10} 0.0$$06_{0.00}^{0.10}$
DeepSeek-V3 $0.98_{0.93}^{1.00}$ $0.90_{0.90}^{0.90}$ $0.74_{0.60}^{0.83}$ $0.62_{0.50}^{0.77}$ $0.49_{0.33}^{0.67}$ $0.49_{0.33}^{0.73}$ $17_{0.13}^{0.23} 0.0$$7_{0.00}^{0.13} 0.0$$2_{0.00}^{0.03} 0.0$$00_{0.00}^{0.00}$
DeepSeek-V3-2503 $1.00_{1.00}^{1.00}$ $0.94_{0.90}^{0.97}$ $0.96_{0.87}^{1.00}$ $0.83_{0.80}^{0.87}$ $0.70_{0.67}^{0.77}$ $0.66_{0.57}^{0.70}$ $0.39_{0.27}^{0.47}$ $10_{0.10}^{0.10} 0.0$$1_{0.00}^{0.03} 0.0$$01_{0.00}^{0.03}$
DeepSeek-R1 $1.00_{1.00}^{1.00}$ $0.99_{0.97}^{1.00}$ $7_{0.97}^{0.97} 0.990$$9_{0.97}^{1.00}$ $0.87_{0.87}^{0.87}$ $0.78_{0.77}^{0.80}$ $0.62_{0.57}^{0.67}$ $24_{0.20}^{0.30} 0.0$$3_{0.00}^{0.10} 0.0$$00_{0.00}^{0.00}$
o1-mini $0.84_{0.73}^{0.90}$ $0.89_{0.87}^{0.93}$ $0.67_{0.60}^{0.77}$ $0.57_{0.43}^{0.63}$ $0.34_{0.23}^{0.47}$ $0.37_{0.23}^{0.43}$ $18_{0.07}^{0.30} 0.0$$1_{0.00}^{0.03} 0.0$$_{0.00}^{0.00} 0.000$$0_{0.00}^{0.00}$
o3-mini $0.79_{0.73}^{0.87}$ $0.62_{0.53}^{0.67}$ $0.53_{0.47}^{0.63}$ $0.28_{0.20}^{0.33}$ $0.31_{0.23}^{0.37}$ $0.30_{0.17}^{0.47}$ $0.30_{0.20}^{0.37}$ $0.19_{0.17}^{0.20}$ $12_{0.07}^{0.17} 0.0$$07_{0.00}^{0.13}$